# Task-free Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

### v3

@@ 4.2.1 Learn from Errors by Selective Weighting
If a trial fails, we compare each image in the retrieved demonstrations against all images from the failed trajectories using $L_{2}$ distance of their embeddings. This comparison yields an Embedding Distance Matrix (EDM, shown in Figure 10) for each retrieved demonstration, where each value represents an embedding distance of a demonstration frame and an image from the failed rollout. This metric determines whether a particular frame has occurred during a failed rollout. Through this process, we identify the Separation Segment — frames in a demonstration where the failed rollout’s behavior starts to diverge from the demonstration (see Figure 4). Since these Separation Segments highlight expected behaviors that did not occur, we consider them critical points contributing to failure. We assign higher weights to these frames which will scale the losses during local adaptation. Detailed heuristics and implementation specifics are provided in Appendix A.4.
@@ 4.2.2 Local Adaptation with Fast Finetuning
Finally, we fine-tune the network’s parameters to better adapt to the deploying scenario $S_{deploy}$ using the retrieved demonstrations from $\mathcal{M}$, focusing more on the Separation Segments identified through selective weighting. Despite this limited data, our experiments demonstrate that the model can effectively recover learned knowledge and skills and improve robot’s performance across various tasks. Overall, the proposed weighted local adaptation is formalized as follows:
$\mathbf{\theta}^{*}=\arg\min_{\mathbf{\theta}}\sum_{n=1}^{\tilde{N}}\sum_{t=1}^{l_{n}}w_{t,n}\mathcal{L}\left(\pi_{\theta}(o_{\leq t,n},g_{n}),a_{t,n}\right)$ (2)
where $\tilde{N}$ is the number of retrieved demonstrations, $l_{n}$ is the length of demonstration $n$, and $w_{t,n}$ is the weight assigned to sample $t$ in demonstration $n$. The variables $o_{\leq t,n}$ and $a_{t,n}$ denote the sequence of observations up to time $t$ and the corresponding expert action, respectively, while $g_{n}$ is the language description for demonstration $n$. The parameter $\mathbf{\theta}$ represents the network’s parameters before local adaptation.
@@ 5 Experiments
We conduct a comprehensive set of experiments to evaluate the effectiveness of RWLA for task-free lifelong robot learning. Specifically, our experiments aim to address the following key questions:
@@ 5.1.1 Benchmarks
We paraphrased the assigned single task goal into diverse language descriptions to obscure task boundaries (See Figure 7). These enriched language descriptions were generated by rephrasing the original task goal from the benchmark using a large language model provided by Phi-3-mini-4k-instruct Model (mini-4k instruct, 2024), ensuring consistent meanings while varying phraseology and syntax. Please see Section A.3 for more details.
@@ 5.1.2 Baselines
Table 1: Comparison with Baselines. The Average Success Rates (ASR, %) and standard deviations (STD, %) across various baselines are shown below. We provide PackNet’s performance as a reference point for cases where task IDs are accessible. Both EWC and vanilla AGEM demonstrate weak performance across all benchmarks. Under our Retrieval-based Weighted Local Adaptation (RWLA) paradigm, both ER and AGEM show significant improvements over their vanilla counterparts, highlighting the effectiveness of RWLA.
@@
EWC PackNet AGEM AGEM-RWLA ER ER-RWLA (ours)
libero_spatial 0.0 $\pm$ 0.0 53.17 $\pm$ 7.25 7.33 $\pm$ 1.61 35.83 $\pm$ 9.39 15.67 $\pm$ 6.66 39.83 $\pm$ 8.13
libero_object 1.50 $\pm$ 0.50 73.67 $\pm$ 1.89 27.17 $\pm$ 4.65 51.17 $\pm$ 3.62 56.50 $\pm$ 3.97 62.33 $\pm$ 1.76
libero_goal 0.33 $\pm$ 0.58 66.33 $\pm$ 4.19 10.83 $\pm$ 3.55 58.67 $\pm$ 3.21 52.33 $\pm$ 5.20 62.33 $\pm$ 4.93
libero_different_scenes 2.58 $\pm$ 1.14 32.92 $\pm$ 0.63 20.43 $\pm$ 1.17 41.75 $\pm$ 5.66 34.08 $\pm$ 2.47 45.17 $\pm$ 0.38
Table 2: Ablation Study on Selective Weighting. This table presents ASR (%) and their STD (%) for uniform (RULA) and weighted (RWLA) local adaptation across $15$, $20$, and $25$ epochs of adaptation under three random seeds, with evaluations conducted on all $10$ tasks within the benchmarks: libero_spatial, libero_object, and libero_goal. Compared to RULA, selective weighting scheme improves the method’s performance on most benchmarks.
Benchmark Method 15 Epochs 20 Epochs 25 Epochs Overall ASR
libero_spatial RULA 35.33 $\pm$ 9.07 38.17 $\pm$ 5.06 38.17 $\pm$ 3.33 37.22 $\pm$ 5.64
RWLA 36.17 $\pm$ 8.33 39.83 $\pm$ 8.13 37.83 $\pm$ 3.75 37.94 $\pm$ 6.32
libero_object RULA 57.83 $\pm$ 9.07 60.67 $\pm$ 3.01 58.00 $\pm$ 4.92 58.83 $\pm$ 5.55
RWLA 58.00 $\pm$ 4.44 62.33 $\pm$ 1.76 61.50 $\pm$ 1.32 60.61 $\pm$ 3.18
libero_goal RULA 61.33 $\pm$ 3.69 62.00 $\pm$ 2.60 66.17 $\pm$ 5.58 63.17 $\pm$ 4.24
RWLA 62.83 $\pm$ 7.65 62.33 $\pm$ 4.93 67.50 $\pm$ 1.80 64.22 $\pm$ 5.26
@@ 5.1.4 Model, Training, and Evaluation
As illustrated in Figure 2, our model utilizes pretrained encoders for visual and language inputs: R3M (Nair et al., 2022) for visual encoding, Sentence Similarity model (SS Model) (SentenceSimilarity, 2024) for language embeddings, and a trainable MLP-based network to encode proprioceptive inputs. Embeddings from ten consecutive time steps are processed through a transformer-based temporal encoder, with the resulting output passed to a GMM-based policy head for action sampling. Specifically, R3M, a ResNet-based model trained on egocentric videos using contrastive learning, captures temporal dynamics and semantic features from scenes, while Sentence Similarity Model captures semantic meanings in language descriptions, enabling the model to differentiate between various instructions.
@@
After training on all tasks sequentially, we conduct reviewing and testing on various scenarios sampled from each task for comprehensive analysis. During the reviewing stage, we firstly evaluate potential forgetting by having the agent perform $10$ rollout episodes on the deployment scenario $\mathcal{S}_{deploy}$. We then retrieve the most similar demonstrations from $\mathcal{M}$ and fine-tune the model for only $20$ epochs using the retrieved demonstrations with selective weighting. Finally, we deploy the adapted model for $20$ episodes—the testing phase—to assess performance improvements. All training, local adaptation, and testing in the benchmarks are conducted using three random seeds ($1$, $21$, and $42$) to reduce the impact of randomness.
@@ 5.2.1 Comparison with Baselines
In contrast, PackNet serves as a reference method, as it requires explicit task IDs. Noteworthy, as the number of tasks increases, the network’s trainable capacity under PackNet diminishes, leaving less flexibility for future tasks. This limitation becomes evident in the libero_different_scenes benchmark, which includes 20 tasks, see Table 8. PackNet’s success rate drops significantly for later tasks, resulting in poor overall performance and highlighting its constraints on plasticity compared with the proposed ER-RWLA.
@@
(a) PCA based on BERT
(b) PCA based on SS Model
(c) Bar Chart of Average Success Rates and Retrieval Accuracy across $10$ tasks
Figure 5: In Figures 5(a) and 5(b), PCA is used to visualize the distribution of language embeddings of 3 tasks from BERT and SS, respectively. In Figure 5(c), the SS model, which distinguishes task descriptions, has higher average success rates (ASR) and retrieval accuracy (RA) than BERT. The error bars represent the standard deviations of ASR and RA for each task over 20 repetitions with 3 random seeds.
@@ 5.2.2 Ablation Studies
In the first ablation, we evaluated the impact of selective weighting on libero_spatial, libero_object, and libero_goal benchmarks to demonstrate its importance for effective local adaptation. We compared a variant of RWLA: RULA, which applies uniform local adaptation without selective weighting, adapting retrieved demonstrations uniformly. Both methods are trained with ER.
Since early stopping during local adaptation at test time is infeasible, and training can be unstable, particularly regarding manipulation success rates, we conducted RWLA using three different numbers of epochs — $15$, $20$, and $25$. The results presented in Table 2, indicate that selective weighting enhances performance across different adaptation durations and various benchmarks, addressing Question 3.
@@
(2) Due to this limitation, BERT is unable to retrieve the most relevant demonstrations (those most similar to the current scenario from the storage memory $\mathcal{M}$). As a result, RWLA with BERT does not achieve optimal performance. These two findings address Question 1.
(3) Interestingly, from Figure 5 (c), despite BERT’s low Retrieval Accuracy (RA), if it attains a moderately acceptable rate (e.g., 0.375), the RWLA based on BERT embeddings can still enhance model performance. This demonstrates the robustness and fault tolerance of our proposed approach, further addressing Question 4 and 5.
@@ 6 Conclusion and Discussion
In this paper, we introduced a novel task-free lifelong robot learning framework that combines retrieval-based local adaptation with selective weighting before policy deployment. Our approach enables robots to continuously learn and restore forgotten knowledge and skills in dynamic environments without relying on task boundaries or IDs. Notably, our framework is not only robust, but is compatible with various memory-based lifelong learning methods, enhancing a robot’s ability to perform previously learned tasks as a paradigm.
@@
Appendix A Appendix
@@ A.1 Notations
$\mathcal{M}$ Episodic memory buffer
$\mathbf{o}_{t}$ Observation vector at time $t$
$\mathbf{o}_{\leq t}$ Sequence of observation vectors up to time $t$ to deal with partial observability
$\mathbf{a}_{t}$ Action vector at time $t$
$\mathbf{a}_{t}^{k}$ Action vector at time $t$ for task $k$
@@
$\mathbf{\theta}_{k}$ Model parameters after adaptation on task $k$
@@
$\pi_{\theta}(\mathbf{s}_{\leq t},\mathcal{T}_{k})$ Policy output given states up to time $t$ and task $\mathcal{T}_{k}$
@@
Algorithm 1 RWLA for Task-free Lifelong Robot Learning
Lifelong Learning Phase:
  1. 1.
  2. 2.
    $K\in\{1,2,\dots,T\}$
    1. (a)
      Train $\theta$ on $\mathcal{D}_{K}\cup\mathcal{M}$ using Eq 1
    2. (b)
Reviewing Phase:
  1. 1.
  2. 2.
    Retrieve $\tilde{N}$ demonstrations from $\mathcal{M}$ based on embedding distance using Eq 4.1 (4.1)
  3. 3.
    Compute $w_{t,n}$ based on selective weighting (4.2.1)
  4. 4.
    $\theta^{\prime}\leftarrow$ Locally adapt $\theta$ using Eq 2 as skill restoration within limited epochs (4.2.2)
Testing Phase: Test $\theta^{\prime}$ in $\mathcal{S}_{deploy}$
@@ A.2.1 Network Architecture and Modularities
Pretrained Language Encoder Sentence Similarity (SS) Model (SentenceSimilarity, 2024), output size: 384
@@
Temporal Transformer heads: 6, sequence length: 10, dropout: 0.1, head output size: 64
@@
Action $7$
@@ A.2.2 Training Hyperparameters
Task performance was evaluated every $10$ epochs using $20$ parallel processes to maximize efficiency. The best-performing model from these evaluations was retained for subsequent tasks. After training on each task, we reassessed the model’s performance across all previously encountered tasks.
@@
Number of Demos per Task in $\mathcal{M}$ ($\tilde{N}$) $8$
@@
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_spatial and libero_object $[1.0,0.5]$
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_goal $[0.5,1.0]$
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_different_scenes $[1.0,0.1]$
@@
Clipping Range for Selective Weighting $2$
Default Local Adaptation Epochs $20$
@@ A.3.1 Datasets Structure
Our dataset inherent the dataset from LIBERO (Liu et al., 2024), maintaining all the attributes and data. Additionally, we add demo description to each demonstration to blur task boundary and augment language description during training (See Figure 6). Unlike the dataset from LIBERO, which groups demonstrations together under one specific task, our dataset wrap all demonstrations with random order to eliminate the task boundary.

### v4

@@ 4.2 Weighted Local Adaptation
Figure 3: Storage memory $\mathcal{M}$ retains a small set of demonstrations during lifelong learning. To retrieve those most relevant to the current scenario, we compute a weighted distance (Eq 2) using image and language embeddings. Red and green circles indicate relevant and irrelevant demonstrations, respectively, which include language, visual observations, expert actions, and robot states. Retrieved demonstrations are then used for Weighted Local Adaptation.
@@ 4.2.1 Learn from Errors by Selective Weighting
If a trial fails, we compare each image in the retrieved demonstrations against all images from the failed trajectories using $L_{2}$ distance of their embeddings. This comparison yields an Embedding Distance Matrix (EDM, shown in Figure 11) for each retrieved demonstration, where each value represents an embedding distance of a demonstration frame and an image from the failed rollout. This metric determines whether a particular frame has occurred during a failed rollout. Through this process, we identify the Separation Segment — frames in a demonstration where the failed rollout’s behavior starts to diverge from the demonstration (see Figure 4). Since these Separation Segments highlight expected behaviors that did not occur, we consider them critical points contributing to failure. We assign higher weights to these frames, which will scale the losses during local adaptation. Detailed heuristics and implementation specifics are provided in Appendix A.4.
@@ 4.2.2 Local Adaptation with Fast Finetuning
Finally, we fine-tune the network’s parameters to better adapt to the deployment scenario $S_{deploy}$ using the retrieved demonstrations from $\mathcal{M}$, focusing more on the Separation Segments identified through selective weighting. Despite this limited data, our experiments demonstrate that the model can effectively recover learned knowledge and skills and improve the robot’s performance across various tasks. Overall, the proposed weighted local adaptation is formalized as follows:
$\theta^{*}=\arg\min_{\theta}\sum_{n=1}^{\tilde{N}}\sum_{t=1}^{l_{n}}w_{t,n}\mathcal{L}\left(\pi_{\theta}(o_{\leq t,n},g_{n}),a_{t,n}\right),$ (3)
where $\tilde{N}$ is the number of retrieved demonstrations, $l_{n}$ is the length of demonstration $n$, and $w_{t,n}$ is the weight assigned to sample $t$ in demonstration $n$. The variables $o_{\leq t,n}$ and $a_{t,n}$ denote the sequence of observations up to time $t$ and the corresponding expert action, respectively, while $g_{n}$ is the language description for demonstration $n$. The parameter $\theta$ represents the network’s parameters before local adaptation.
@@ 5 Experiments
We conduct a comprehensive set of experiments to evaluate the effectiveness of RWLA for task-agnostic lifelong robot learning. Specifically, our experiments aim to address the following key questions:
@@ 5.1.1 Benchmarks
We paraphrased the assigned single task goal into diverse language descriptions to obscure task boundaries (See Figure 7). These enriched language descriptions were generated by rephrasing the original task goal from the benchmark using a large language model provided by Phi-3-mini-4k-instruct Model (mini-4k instruct, 2024), ensuring consistent meanings while varying phraseology and syntax. Please see Appendix A.3 for more details.
@@ 5.1.2 Baselines
Table 1: Comparison with Baselines. The Average Success Rates (ASR, %) across various baselines are shown below. We provide PackNet’s performance as a reference point for cases where task IDs are accessible. Both EWC and vanilla AGEM demonstrate weak performance across all benchmarks. Under our Retrieval-based Weighted Local Adaptation (RWLA) paradigm, both ER and AGEM show significant improvements over their vanilla counterparts, highlighting the effectiveness of RWLA.
@@
[…]
@@ 5.1.4 Model, Training, and Evaluation
As illustrated in Figure 2, our model utilizes pretrained encoders for visual and language inputs: R3M (Nair et al., 2022) for visual encoding, Sentence Similarity model (SS Model) (SentenceSimilarity, 2024) for language embeddings, and a trainable MLP-based network to encode proprioceptive inputs. Embeddings from ten consecutive time steps are processed through a transformer-based temporal encoder, with the resulting output passed to a GMM-based policy head for action sampling. Specifically, R3M, a ResNet-based model trained on egocentric videos using contrastive learning, captures temporal dynamics and semantic features from scenes, while the Sentence Similarity Model captures semantic meanings in language descriptions, enabling the model to differentiate between various instructions.
@@
After training on all tasks sequentially, we conduct reviewing and testing on various scenarios sampled from each task for comprehensive analysis. During the reviewing stage, we firstly evaluate potential forgetting by having the agent perform $10$ rollout episodes on the deployment scenario $\mathcal{S}_{deploy}$. We then retrieve the most similar demonstrations from $\mathcal{M}$ and fine-tune the model for only $20$ epochs using the retrieved demonstrations with selective weighting. Finally, we deploy the adapted model for $20$ episodes—the testing phase—to assess performance improvements. All training, local adaptation, and testing in the benchmarks are conducted using three random seeds ($1$, 2121, and $42$) to reduce the impact of randomness.
@@ 5.2.1 Comparison with Baselines
In contrast, PackNet — relying on explicit task IDs — allocates a separate slice of network weights to each new task. Noteworthy, this strategy works well early on, but as the number of tasks increases, the network’s trainable capacity under PackNet diminishes, leaving less flexibility for future tasks. This limitation becomes evident in the libero_different_scenes benchmark, which includes 20 tasks, see Appendix A.6. PackNet’s success rate drops significantly for later tasks, resulting in poor overall performance and highlighting its constraints on plasticity compared with the proposed ER-RWLA.
@@
Table 2: Ablation Study on Selective Weighting. This table presents ASR (%) for uniform (RULA) and weighted (RWLA) local adaptation across $15$, 2020, and $25$ epochs of adaptation under three random seeds, with evaluations conducted on all $10$ tasks within the benchmarks: libero_spatial, libero_object, and libero_goal. Compared to RULA, selective weighting scheme improves the method’s performance on most benchmarks.
Benchmark Method 15 Epochs 20 Epochs 25 Epochs Overall ASR
libero_spatial RULA 35.33 38.17 38.17 37.22
RWLA 36.17 39.83 37.83 37.94
libero_object RULA 57.83 60.67 58.00 58.83
RWLA 58.00 62.33 61.50 60.61
libero_goal RULA 61.33 62.00 66.17 63.17
RWLA 62.83 62.33 67.50 64.22
@@ 5.2.2 Ablation Studies
In the first ablation, we evaluated the impact of selective weighting on libero_spatial, libero_object, and libero_goal benchmarks to demonstrate its importance for effective local adaptation. We compared a variant of RWLA: RULA, which applies uniform local adaptation without selective weighting, adapting retrieved demonstrations uniformly. Both methods are trained with ER. Since early stopping during local adaptation at test time is infeasible, and training can be unstable, particularly regarding manipulation success rates, we conducted RWLA using three different numbers of epochs — $15$, 2020, and $25$.
The results presented in Table 2 indicate that selective weighting enhances performance across different adaptation durations and various benchmarks, addressing Question 3. The gains over uniform adaptation are modest but systematic, echoing findings from (Byrd & Lipton, 2019), that importance weighting does not yield substantial performance jumps once expressive models have converged. Crucially, RWLA achieves these consistent improvements with negligible overhead — only a handful of demonstrations and a few extra adaptation epochs — making selective weighting a practical, low-cost boost for reliable skill recovery.
@@
(a) PCA based on BERT
(b) PCA based on SS Model
(c) Bar Chart of Average Success Rates (ASR) and Retrieval Accuracy (RA) across $10$ tasks
Figure 5: In Figures 5(a) and 5(b), PCA is used to visualize the distribution of language embeddings of 3 tasks from BERT and $S$, respectively. In Figure 5(c), the SS model, which distinguishes task descriptions, has higher average success rates (ASR) and retrieval accuracy (RA) than BERT. The error bars represent the standard deviations of ASR and RA for each task over 20 repetitions with 3 random seeds.
@@
(2) Due to this limitation, BERT is unable to retrieve the most relevant demonstrations (those most similar to the current scenario from the storage memory $\mathcal{M}$). As a result, RWLA with BERT does not achieve good performance. These two findings address Question 1.
(3) Interestingly, from Figure 5 (c), despite BERT’s low Retrieval Accuracy (RA = proportion of the top $\tilde{N}$ demonstrations retrieved from $\mathcal{M}$ that come from the ground-truth task of $\mathcal{S}_{deploy}$), if it attains a moderately acceptable rate (e.g., 0.375), the RWLA based on BERT embeddings can still enhance model performance. This demonstrates the robustness and fault tolerance of our proposed approach, further addressing Questions 4 and 5.
@@ 6 Conclusion and Discussion
We introduce a task-agnostic lifelong robot learning framework that combines retrieval-based local adaptation with selective weighting before policy deployment. During continual learning, when performance on a previously learned task degrades, the robot triggers an on-demand “review” phase that uses just a few stored demonstrations to rapidly recover the forgotten skills — without requiring task IDs or boundaries. The method is robust and plug-and-play, seamlessly enhancing the memory-based lifelong learning robot’s ability to retain and recover prior knowledge while it continues to acquire new skills.
@@
Appendix A Appendix for Task-agnostic Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation
@@ A.1 Notations
$\mathcal{M}$ Storage memory buffer
$o_{t}$ Observation vector at time $t$
$o_{\leq t}$ Sequence of observation vectors up to time $t$ to deal with partial observability
$a_{t}$ Action vector at time $t$
$a_{t}^{k}$ Action vector at time $t$ for task $k$
@@
$\theta_{k}$ Model parameters after adaptation on task $k$
@@
$\pi_{\theta}(s_{\leq t},\mathcal{T}_{k})$ Policy output given states up to time $t$ and task $\mathcal{T}_{k}$
@@ A.2.1 Network Architecture and Modularities
Pretrained Language Encoder
Sentence Similarity (SS) Model (SentenceSimilarity, 2024),
output size: 384
@@
Temporal Transformer
heads: 6, sequence length: 10,
dropout: 0.1, head output size: 64
@@
Action 77
Image Encoding: We employ the R3M visual encoder (Nair et al., 2022), pretrained on large-scale robotic data, to extract rich image features. These features are then passed through a lightweight MLP that projects them into a compact 64-dimensional embedding.
Language Encoding: Natural language instructions are first embedded by a sentence-similarity model (SentenceSimilarity, 2024), yielding a 384-dimensional vector. A lightweight MLP then compresses this into a 64-dimensional latent representation, aligned with the other modalities.
Proprioceptive Encoding: Joint angles and gripper states are processed through a small MLP stack, whose final layer also outputs a 64-dimensional vector. This ensures that vision, language, and proprioception share the same embedding size.
Temporal Transformer: At each timestep, we concatenate the three 64-dim modality embeddings and feed the sequence (up to 10 steps) into a 4-layer Transformer decoder. Each layer has 6 attention heads (64-dim head outputs), a 256-unit MLP, sinusoidal positional encodings, and 0.1 dropout to guard against overfitting.
Policy Head: The Transformer’s final output is passed to a GMM-based head with 2 fully-connected layers. It predicts a 5-component Gaussian mixture—outputting per-mode means, standard deviations (clamped $\geq$ 1e-4), and mixture logits.
@@ A.2.2 Training Hyperparameters
Task performance was evaluated every $10$ epochs using $20$ parallel processes to maximize efficiency. The best-performing model from these evaluations was retained for subsequent tasks. After training on each task, we reassessed the model’s performance across all previously encountered tasks. Although we report results only at the end of the full sequence, RWLA remains fully compatible with intermittent skill recovery during continuous deployment: it can be invoked when performance degradation happens, mirroring realistic lifelong-learning robots that adapt on an as-needed basis.
@@
Number of Demos per Task in $\mathcal{M}$ ($\tilde{N}$) 88
@@
Distance weights $\alpha_{v},\alpha_{l}]$ for libero_spatial and libero_object $[1.0,0.5]$
Distance weights $\alpha_{v},\alpha_{l}]$ for libero_goal $[0.5,1.0]$
Distance weights $\alpha_{v},\alpha_{l}]$ for libero_different_scenes $[1.0,0.1]$
@@
Clipping Range for Selective Weighting 22
Default Local Adaptation Epochs 2020
@@ A.3.1 Datasets Structure
Our dataset inherent the dataset from LIBERO Liu et al. (2024), maintaining all the attributes and data. Additionally, we add demo description to each demonstration to blur task boundary and augment language description during training (See Figure 6). Unlike the dataset from LIBERO, which groups demonstrations together under one specific task, our dataset wrap all demonstrations with random order to eliminate the task boundary.